OpenStudy (anonymous):

Please can you help? A poker hand (5 cards) is dealt. i Find the chance that the cards are all diamond. My Solution approach: i)P(all diamond)= (13/52)*(12/51)*(11/50)*(10/49)*(9/48)= yet its wrong????

OpenStudy (agent0smith):

Use combinations instead: From 13 diamonds, choose 5. Divide this by from 52 cards choose 5. (13C5)/(52C5)

OpenStudy (anonymous):

That is indeed the correct answer. Are there more then 1 player(s) though?

OpenStudy (anonymous):

agent0smith and your solutions are in fact equivalent

OpenStudy (anonymous):

can You please in percentages or numbers can you quantify 13C5/52C5=

OpenStudy (agent0smith):

\[\Large nCr =\frac{ n! }{ (n-r)! r! }\] 13C5 = 1287 52C5 = 2598960

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You guys had the exact same answer!!! just written differently. Its important for you to know that.

OpenStudy (agent0smith):

Yeah, I'm guessing you entered it wrongly on your calculator, if it was wrong.